You deposit $12,000 dollars in an account that pays 5.5% interest compounded monthly. What would your balance be after two years?
What if the deposit in above question was compounded continuously? What would the balance be after 2 years?

Answer by fractalier(6550)

Did this one yesterday...see my earlier solution...
$13391.86
Continuously, we have
A(2) = Pe^rt = 12000*e^(.055 * 2) =
$13395.34